# Towards Stealthy and Effective Backdoor Attacks on Lane Detection: A Naturalistic Data Poisoning Approach

### v1

@@ Introduction
RQ4: Physical Experiment
To better simulate real-world conditions, we go beyond prior sandbox setups [16] and conduct physical tests in four closed-road environments. We use two types of cones with different shapes and colors to simulate diverse trigger appearances and evaluate attack performance on the infected LaneATT model under the LOA setting.
We collect 100 poisoned images using cones across both daytime and nighttime conditions with diverse environments and placements, yielding a 57% ASR. The findings suggest that DBALD remains potentially effective under varying conditions in physical scenarios, and thus poses a real threat to LD models in the physical world. Note that, although lower than simulation-based results, it is expected as the model is not retrained on the physical domain. Additionally, using SOTA weather transfer methods [59, 60], we apply rain, blur, and occlusion to 100 poisoned images, observing only slight ASR drops (from 57% to 45/48/49%), indicating our method’s robustness to diverse environments. Figure 7 shows benign and attack cases across different cones and lighting conditions. More results can be viewed on [53].
Figure 7: Physical-world attacks under different conditions.
Countermeasures
Table 7: Defense results (%) of neuron pruning.
Num 0 50 100 150 200 250 300 350 400 450
ACC 74.20 73.70 73.53 72.31 69.26 60.04 53.74 43.97 21.19 3.16
ASR 74.19 73.67 73.37 69.34 62.41 51.84 42.60 34.46 15.87 1.78
To evaluate DBALD against potential backdoor defenses, following existing work [16], we employ fine-tuning and neural pruning as defense strategies. Considering LaneATT under LOA, we fine-tune the attacked LD model on clean samples in 50 epochs, and after 50 epochs, the ASR decreases from 74.19% to 63.68%. This indicates that fine-tuning indeed somewhat mitigates the attack from DBALD but fails to eliminate it. For neural pruning, we prune the last convolutional layer in the model backbone, which contains 512 neurons, starting from 0 with a step size of 50. Results are shown in Table 7. However, we find that pruning a small number of neurons has little impact on the backdoor attack, while pruning more neurons accelerates performance degradation on clean samples. This suggests that neuron pruning is largely ineffective as a defense strategy against our attack. We also assess a vision-language model Qwen2.5-VL[61] as an anomaly detector. Sampling 100 poisoned images, we use the model to identify visual anomalies. Detection rates are low: only 6% for cone-based triggers and 3% for mud-based triggers, indicating that multimodal models struggle to detect our triggers.
Conclusion
We propose DBALD, a diffusion-based backdoor framework for LD systems that places triggers in sensitive regions guided by gradient attention. It supports multiple attack strategies via strategy-specific heatmaps and improves stealth through lane preservation and environment consistency losses. Experiments show DBALD outperforms prior backdoors in both digital and physical settings. Regarding future improvements, we believe a task-specific distilled diffusion model could enhance efficiency. While we currently use UltraEdit as a general-purpose backbone, a tailored model trained to seamlessly embed triggers in driving scenes may substantially reduce diffusion steps and costs.

### v2

@@ Appendix B More Experiment results
Justification of the attention mechanism
The proposed gradient-based attention mechanism is grounded in both prior work and our analytical findings.
Prior work [wenger2021backdoor] demonstrates that gradient-based high-attention regions are strongly correlated with high ASR, suggesting that effective triggers must be strategically placed. Motivated by this, we introduce a similar gradient-driven strategy tailored to lane detection (LD) tasks.
We further conduct an analytical study by visualizing the gradient-based attention maps over training epochs for both clean and trigger-injected images. We observe that:
  • For clean images, attention gradually concentrates on task-relevant semantic regions.
  • For trigger-injected images, attention becomes increasingly focused on the trigger region across training.
This evolving focus indicates that the model gradually “memorizes” the trigger position, reinforcing the effectiveness of targeted trigger placement. Table 6 summarizes results at epoch 100.
Table 6: Comparison of gradient-based attention maps at epoch 100.
Setting Attention Entropy $\downarrow$ Attention on Trigger Region $\uparrow$
Trigger (Ours) 16.09 47.39%
Random Trigger 16.26 44.13%
Clean Image 16.96 41.58%
The lower entropy and higher focus on the trigger region for our method suggest that the model allocates more stable and intense attention to the trigger, thereby improving attack reliability.
Computation time
DBALD incurs a moderate computational cost relative to baseline methods (6 seconds per image vs. 1 second). Since data poisoning is a one-time offline procedure, this computational overhead is generally acceptable.
Clarification on trigger diversity
To preserve high clean accuracy and maintain a low false-positive (FP) rate—particularly in the rare cases where benign patterns resemble potential triggers—we use LPIPS scores to distinguish our triggers from benign visual patterns in the original data. We adopt a threshold of 0.15 to differentiate benign patterns from actual triggers. On benign-only datasets (e.g., TuSimple), this yields a false-positive rate as low as 8.2%. It is important to note that: i) false positives are inevitable in any backdoor attack when considered in an open-world usage setting; and ii) backdoor triggers can be non-fixed and may even be designed to vary dynamically during inference to achieve high ASR.
In this work, our priority is to achieve high ASRs with realistic, context-aware triggers while keeping the false-positive rate low. To this end, we maintain two lightweight reference databases throughout training and evaluation:
  • Benign pattern database: a collection of naturally occurring benign visual patterns in clean images that are semantically similar to potential triggers.
  • Poisoned trigger database: a repository of all synthesized triggers used during the backdoor injection process.
These two databases ensure that each newly synthesized trigger remains sufficiently distinct from benign patterns (LPIPS >0.15>0.15 when compared against the benign pattern database) while remaining visually consistent with existing poisoned triggers (LPIPS <0.15<0.15 when compared against the poisoned trigger database), thereby preserving clean accuracy, maintaining low false-positive rates, and sustaining high ASR with context-aware triggers.
We note that although the clean dataset is large, benign trigger-like patterns are extremely rare (e.g., mud-like textures do not appear, and cone-like shapes occur in only 129 images), resulting in a very small benign-pattern database; the poison-trigger database is similarly small, containing only about 1% of the data corresponding solely to injected trigger instances. Owing to their small size, the pairwise LPIPS comparisons used for trigger-diversity enforcement incur only millisecond-level overhead and thus do not affect overall training efficiency.
Finally, the results in this paper are demonstrated using only two specific trigger instances. Our framework is inherently designed to construct general, natural, and context-aware triggers, and the LPIPS-based filtering together with the dual-database mechanism generalizes uniformly across diverse trigger types, not limited to mud or cones.
Appendix C More details of LD models
Lane detection (LD) methods can be broadly categorized into two main paradigms: anchor-based and segmentation-based approaches. In this section, we first introduce the core characteristics of each category, followed by detailed descriptions of the four representative models evaluated in our experiments. For fair comparison, we adopt ResNet-34 [he2016deep] as the unified backbone for all models.
Anchor-based Methods
Anchor-based approaches formulate lane detection as an anchor classification and refinement task. These methods rely on predefined geometric priors (e.g., horizontal anchors or curve parameters) to guide lane instance generation. Their architectures typically consist of two main components:
Classification Head: Generates candidate lane anchors and classifies whether each anchor corresponds to a valid lane instance based on confidence scores.
Anchor Refinement Module: Performs fine-grained regression to refine the initial anchor positions and improve coordinate-level prediction accuracy.
LaneATT [tabelini2021keep] adopts horizontal line anchors as structural priors and introduces an attention mechanism to dynamically aggregate features from neighboring anchors, which enhances the model’s perception of long-range lane lines. The model consists of a classification head and a regression head: the classification head determines whether each anchor point belongs to a lane line, optimized using a focal loss; while the regression head estimates the spatial offsets to recover the lane shape, trained with an L1 loss.
ADNet [xiao2023adnet] predicts anchors by decoupling them into starting points and directions via heatmaps, enabling flexible and high-quality anchor generation across the entire image. The classification of lane anchors is also optimized using a focal loss, while the lane shape regression adopts a General Lane IoU (GLIoU) loss to better capture geometric discrepancies in challenging scenarios.
Segmentation-based Methods
Segmentation-based methods treat lane detection as a pixel-wise semantic segmentation task, where the network produces dense lane masks. Post-processing steps, such as clustering or polynomial fitting, are then applied to extract individual lane instances. These models commonly include the following components:
Spatial message-passing mechanisms: enhances contextual interactions between pixels, thereby improving the network’s ability to recover structural information in the presence of occlusions or low-contrast regions.
Segmentation Head: Outputs pixel-level class probabilities and typically incorporates topological constraints to enforce structural consistency of predicted lanes.
SCNN [pan2018spatial] introduces directional spatial convolutions along rows and columns, enabling explicit message propagation within feature maps. The model adopts a categorical cross-entropy (CE) loss for pixel-wise classification: each pixel is assigned to one of several predefined lane classes, allowing the network to learn lane-specific spatial distributions. In parallel, a binary cross-entropy (BCE) loss is applied to supervise lane existence prediction. Although the outputs does not directly display lane coordinates, the predicted probability maps serve as the basis for coordinate extraction during inference. Lane points are sampled by selecting the highest response along each row of the probability map.
RESA [zheng2021resa] proposes a recurrent feature-shifting mechanism that aggregates spatial information across multiple directions by cyclically shifting and updating feature maps. This design allows each pixel to incorporate long-range contextual cues, improving robustness to occlusion and broken lane segments. A BCE loss is employed for pixel-wise classification: each pixel is classified as either foreground (lane) or background. The BCE loss trains the network to produce a binary heatmap in which higher values indicate greater likelihood of lane presence. Although this classification supervision does not directly regress lane coordinates, it provides a dense prediction map from which coordinates are later extracted. During inference, lane points are obtained by sampling the heatmap (e.g., selecting the highest response every few rows). Additionally, a CE loss is used to supervise lane existence prediction through a dedicated classification head.
Table 7: Task-specific Loss Selection for Attack Strategies.
Method Type LDA LOA/LRA
Anchor-based Focal Loss L1 Loss (LaneATT)
GLIoU Loss (AdNet)
Segmentation-based CE Loss BCE Loss
Figure 7: Visualization of our injected mud trigger on the CULane dataset. For each method (including LaneATT, ADNet, RESA, and SCNN), we show the clean input image, the strategy-specific gradient heatmaps used for trigger placement, and the corresponding poisoned images under LDA and LOA/LRA attacks. The highlighted regions indicate high-sensitivity areas where the trigger is placed to maximize the attack effect.
Figure 8: Visualization of the physical backdoor attack in a real-world driving scenario.
Appendix D Heatmap-based Optimal Trigger Position
Building on these foundations, we design a unified gradient-based approach to identify optimal trigger positions for backdoor injection. For a given input image and selected attack strategy—namely Lane Disappearance Attack (LDA), Lane Offset Attack (LOA), or Lane Reconstruction Attack (LRA), we first determine the appropriate loss function according to the model type and task formulation (as summarized in Table 7). Specifically, classification losses such as focal loss or cross-entropy (CE) are used for LDA to suppress lane existence predictions, while regression-oriented losses like L1, BCE, or GLIoU are employed in LOA and LRA to induce geometric distortions. By computing the gradient of the selected loss with respect to the input, we obtain a task-specific heatmap that highlights sensitive regions most influential to the model’s output.
Discussion on Model Transferability Our results show that randomly placed triggers (i.e., without heatmap guidance) can still achieve a related high ASR compared to baselines across all settings. More importantly, we observe a notable degree of cross-model generalization. Triggers generated using the LaneATT heatmap can be transferred to a different lane detection architecture, namely RESA, and still achieve an ASR of 71.49% for the LOA attack. This performance falls between the ASR achieved using RESA-specific heatmaps, which is 76.86%, and random placement, which is 69.37%, suggesting that the heatmap captures model-agnostic vulnerable regions to some extent. Overall, these results indicate that while heatmap guidance improves attack effectiveness, the learned trigger patterns also exhibit a certain level of transferability across different lane detection architectures.
Appendix E More visualizations
Visualization of gradient-based attention heatmaps
Figure 7 presents the visualization of gradient-based attention heatmaps and the resulting physical backdoor trigger placement across different models and attack strategies. For each model, we compute gradient heatmaps corresponding to the LDA and LOA/LRA objectives by leveraging the task-specific loss functions. We then locate the most sensitive regions within the road area by selecting the regions with the highest gradient response values.
As illustrated in the figure, the distribution of gradient attention varies significantly across models. Furthermore, we observe noticeable differences between the LDA and LOA/LRA strategies within the same model. This supports our use of strategy-specific heatmaps for optimizing trigger placement. Notably, the generated mud trigger exhibits an amorphous pattern, which enhances its stealthiness and robustness in physical settings. Such irregular and adaptive shape designs increase the likelihood of successful trigger activation during inference, even under viewpoint or lighting variations.
More physical experiment results
Figure 8 shows real-world evaluations of the physical backdoor attack across four real-world driving scenarios discussed in the main paper. In each scenario, we deploy a cone trigger at different positions along the driving path. Despite the variations in background, lighting, and occlusions (e.g., pedestrians and vehicles), the trigger consistently induces the intended lane detection failures. Notably, the attack remains effective across diverse angles and environmental conditions. These results demonstrate that the LD infected by DBALD can be reliably activated during inference in physical settings. This further confirms that DBALD poses a tangible threat to lane detection models deployed in real-world driving systems.
We also collected 50 real-world samples of the mud attack for physical evaluation. The results demonstrate an ASR of 62%. The trigger visualizations are on [dbald_site].